# Torsion of $\mathbb{Q}$-curves over number fields of small odd prime degree

### v1

@@ Abstract.
We determine all groups which occur as torsion subgroups of $\mathbb{Q}$-curves defined over number fields of degrees $3$, $5$ and $7$. In particular, we prove that every torsion subgroup of a $\mathbb{Q}$-curve defined over a number field of degree $3,5$ or $7$ already occurs as a torsion subgroup of an elliptic curve with rational $j$-invariant. As the quadratic case has been solved by Le Fourn and Najman, and the case of extensions of prime degree greater than $7$ has been solved by Cremona and Najman, this paper completes the classification of torsion of $\mathbb{Q}$-curves over number fields of prime degree. To solve our problem, we describe how images of Galois representations of elliptic curves change upon applying an isogeny. Since Cremona and Najman proved that $\mathbb{Q}$-curves over an odd degree number field $K$ are exactly those which are $K$-isogenous to an elliptic curve with rational $j$-invariant, we can rely on known results about images of Galois representations of rational elliptic curves.
As a corollary of our results, we establish that the torsion subgroup an elliptic curve over a number field $K$ of prime degree which is isogenous to an elliptic curve with rational $j$-invariant is equal to the torsion subgroup of some elliptic curve defined over a degree $p$ number field with rational $j$-invariant.
@@ 1. Introduction
Let $E/K$ be an elliptic curve, where $K$ is a number field. By the Mordell-Weil theorem, the group $E(K)$ is finitely generated and hence of the form $T\times\mathbb{Z}^{r}$ for some $r\geq 0$, where $T$ is a finite Abelian group.
A typical question one might ask is to determine all possible torsion subgroups occuring in a given class of elliptic curves. For elliptic curves over $\mathbb{Q}$, we have the following famous result by Mazur.
@@
[…]
@@
An elliptic curve $E$ over a number field $K$ is a $\mathbb{Q}$-curve if it is isogenous over $\overline{K}$ to all its Galois conjugates.
For example, all elliptic curves isogenous to an elliptic curve with rational $j$-invariant are $\mathbb{Q}$-curves. Indeed, if $E$ is isogenous to an elliptic curve $E^{\prime}$ with rational $j$-invariant, then there is a twist $E^{\prime\prime}$ of $E^{\prime}$ which is defined over $\mathbb{Q}$, and an isogeny $\varphi:E\to E^{\prime\prime}$. Then there is also an isogeny $\varphi^{\sigma}:E^{\sigma}\to(E^{\prime\prime})^{\sigma}$, but $(E^{\prime\prime})^{\sigma}=E^{\prime\prime}$, and the composition of $\varphi$ and the dual of $\varphi^{\sigma}$ is an isogeny from $E$ to $E^{\sigma}$.
Over number fields of odd degree, Cremona and Najman have proven that all $\mathbb{Q}$-curves arise in this way, with the isogeny even being defined over the base field of the $\mathbb{Q}$-curve.
@@
As any isogeny is a composition of a cyclic isogeny and multiplication by $m$ for some integer $m$, we may assume that the isogeny from the above proposition is cyclic.
The goal of this article is to find all possible groups which arise as torsion subgroups of $\mathbb{Q}$-curves over cubic, quintic and septic extensions of $\mathbb{Q}$.
@@
[16, Theorem 1.] Let $E/\mathbb{Q}$ be a rational elliptic curve, and let $K/\mathbb{Q}$ be a cubic extension. Then $E(K)_{tors}$ is isomorphic to one of the following $20$ groups:
[…]
@@
[6, Theorem 1] Let $E/\mathbb{Q}$ be a rational elliptic curve, and let $K/\mathbb{Q}$ be a quintic extension. Then $E(K)_{tors}$ is isomorphic to one of the following $17$ groups:
[…]
@@
[7, follows from Proposition 7.7] Let $E/\mathbb{Q}$ be a rational elliptic curve, and let $K/\mathbb{Q}$ be a septic extension. Then $E(K)_{tors}$ is isomorphic to one of the $15$ groups from Theorem 1.1.
@@
In fact, Gužvić proved in [9, Theorem 5.1.2] that elliptic curves with rational $j$-invariant over an odd prime degree number field have the same list of possible torsion subgroups as base changes of rational elliptic curves.
Over quadratic number fields, there is one exception. Namely, the group $\mathbb{Z}/13\mathbb{Z}$ appears among torsion subgroups of elliptic curves with rational $j$-invariant, and does not appear among torsion subgroups of rational elliptic curves.
@@
Let $K/\mathbb{Q}$ be a cubic extension and let $E/K$ be a $\mathbb{Q}$-curve. Then $E(K)_{tors}$ is isomorphic to one of the $20$ groups appearing in Theorem 1.4.
@@
Let $K/\mathbb{Q}$ be a quintic extension and let $E/K$ be a $\mathbb{Q}$-curve. Then $E(K)_{tors}$ is isomorphic to one of the $17$ groups appearing in Theorem 1.5.
@@
Let $K/\mathbb{Q}$ be a septic extension and let $E/K$ be a $\mathbb{Q}$-curve. Then $E(K)_{tors}$ is isomorphic to one of the $15$ groups appearing in Theorem 1.6.
@@
[…]
@@
In [4, Theorem 1.3.], Cremona and Najman have shown that when $p>7$ is a prime and $K/\mathbb{Q}$ is an extension of degree $p$, every torsion subgroup $E(K)_{tors}$ of some $\mathbb{Q}$-curve $E/K$ is one of the subgroups already appearing in Theorem 1.1.
@@
Under our definition, CM elliptic curves are also $\mathbb{Q}$-curves. Some authors exclude CM curves from the definition of $\mathbb{Q}$-curves, but we will not do that. In any case, the main results of this paper stay the same regardless of the chosen convention. In [3], Clark, Corn, Rice and Stankewicz classified torsion of CM elliptic curves over number fields of degree at most $13$. They proved that groups which occur as torsion subgroups of CM curves over cubic fields are those from Theorem 1.1 along with $\mathbb{Z}/9\mathbb{Z}$ and $\mathbb{Z}/14\mathbb{Z}$. Over quintic fields, the only new group which appears is $\mathbb{Z}/11\mathbb{Z}$. Over septic fields, no new groups appear. In the rest of this article, we only focus on non-CM elliptic curves.
@@
The paper is structured as follows. In Section 2, we list some general helpful results that will be used throughout the paper. In particular, we explicitly spell out the connection between the images of Galois representations of isogenous elliptic curves.
In Section 3, we prove Theorem 1.8, i.e. we solve the problem for cubic extensions. We do this case by case, eliminating each group which occurs for elliptic curves over cubic extensions and does not occur for rational elliptic curves over cubic extensions. Most of the techniques used in the cubic case are also used in the quintic and septic case, which are then discussed in Section 4. In that case, we do not have a complete list of possible torsion subgroups, but we do have finitely many cases to check as there are finitely many degrees of cyclic isogenies of elliptic curves with rational $j$-invariant. Finally, in Section 5, we conclude the proof of Theorem 1.13.
@@ 2. General results
In this section, we recall some helpful results which will be used throughout the paper.
@@
Let $E$ be an elliptic curve over a number field $K$. Let $E^{\prime}$ be a quadratic twist of $E$ by $d\in K^{\times}$. Let $\psi:G_{K}\to\{-1,1\}$ be the quadratic character defined by $\sigma(\sqrt{d})=\psi(\sqrt{d})\sqrt{d}$. Denote by $\rho_{E,N}$ and $\rho_{E^{\prime},N}$ the mod $N$ Galois representations of $E$ and $E^{\prime}$ respectively. Then, for an appropriate choice of bases of $E[N]$ and $E^{\prime}[N]$, we have
@@
The proof can be found in [19, Lemma 5.1].
Since $\mathbb{Q}$-curves are isogenous to elliptic curves with rational $j$-invariants, it is very useful to describe what happens to the image of Galois representations after applying isogenies.
We start with the following lemma which describes what happens with a point of prime order on the starting elliptic curve.
@@
[2, Lemma 4.2.] Let $E_{1}/F$ be an elliptic curve with a point $P\in E_{1}(F)$ of prime order $p$ and let $\varphi:E_{1}\to E_{2}$ be an isogeny defined over $F$. Then either $E_{2}$ has a point of order $p$ or for some basis of $E_{2}[p]$
$\rho_{E_{2},p}(\sigma)=\begin{bmatrix}\chi_{p}(\sigma)&y\\ 0&1\end{bmatrix},\ \text{for all }\sigma\in G_{F}.$
To avoid referring to bases, we may rephrase Lemma 2.2 as follows.
@@
Let $E_{1}/F$ be an elliptic curve with a point $P\in E_{1}(F)$ of prime order $p$ and let $\varphi:E_{1}\to E_{2}$ be an isogeny defined over $F$. Then every $\alpha\in\rho_{E_{2},p}(G_{F})$ has eigenvalue $1$.
Combining Lemma 2.1 and Lemma 2.2, we obtain the following.
@@
Let $E$ be a $\mathbb{Q}$-curve over an odd degree number field $K$ with a point of order $p$ defined over $K$, where $p$ is a prime number. Let $\phi:E\to E^{\prime}$ be an isogeny, where $E^{\prime}$ is a curve with rational $j$-invariant, and let $E^{\prime\prime}/\mathbb{Q}$ be a quadratic twist of $E^{\prime}.$ Then $E^{\prime\prime}$ has a $p$-isogeny defined over $K$ and any element of $\rho_{E^{\prime\prime},p}(G_{K})$ has eigenvalue $1$ or $-1$.
The following proposition, which strengthens [2, Proposition 4.4.], describes what happens in the composite case.
@@
Let $E_{1}/F$ be an elliptic curve with a point $P$ of order $N$ such that $\langle P\rangle$ is fixed by $G_{F}$. Denote by $H\leq(\mathbb{Z}/N\mathbb{Z})^{\times}$ the subgroup $\{k\in(\mathbb{Z}/N\mathbb{Z})^{\times}\mid\sigma(P)=kP\text{ for some }\sigma\in G_{F}\}$.
Let $\phi:E_{1}\to E_{2}$ be a cyclic isogeny defined over $F$. Let $\alpha$ be the order of $\phi(P)$ in $E_{2}$. Then the image of the mod $N$ representation $\rho_{E_{2},N}(G_{F})$ is conjugate to a subgroup of the group of all matrices of the form
@@
Proof.   Let $\{P,Q\}$ be a basis for $E_{1}[N]$. Denote by $\alpha$ and $\beta$ respectively the order of $\phi(P)$ and $\phi(Q)$. Let $R$ be a generator of $E_{1}[N]\cap\ker\phi$. Then $\alpha P=xR$ and $\beta Q=yR$ for some $x,y$. This means that $\alpha\cdot\frac{y}{\gcd(x,y)}P=\beta\cdot\frac{x}{\gcd(x,y)}Q$, so $N$ divides both $\alpha\cdot\frac{y}{\gcd(x,y)}$ and $\beta\cdot\frac{x}{\gcd(x,y)}$. From here, it follows that $N\mid\operatorname{lcm}(\alpha,\beta)$. Namely, assume that $p^{k}$ divides $N$ and doesn’t divide $\alpha$ nor $\beta$. Then $p$ divides both $\frac{y}{\gcd(x,y)}$ and $\frac{x}{\gcd(x,y)}$, a contradiction.
Now let $\{R,S\}$ be a basis for $E_{2}[N]$, chosen so that $\frac{N}{\alpha}R=\phi(P)$ and $\frac{N}{\beta}S=\phi(Q)$. We now compute the action of $\sigma\in G_{F}$ on both elements of this basis.
We know that
@@
Furthermore, $\frac{N}{\beta}\sigma(S)=\sigma(\phi(Q))=\phi(\sigma(Q))$. If we set $\sigma(Q)=uP+h^{-1}\chi_{N}(\sigma)Q$, we get
$\frac{N}{\beta}\sigma(S)=u\phi(P)+h^{-1}\chi_{N}(\sigma)\phi(Q)=\frac{uN}{\alpha}R+h^{-1}\frac{\chi_{N}(\sigma)N}{\beta}S.$
Since the left hand side belongs to $E_{1}[\beta]$, so does the right hand side. Thus, $\frac{N}{\beta}\mid\frac{uN}{\alpha}$, or equivalently $\alpha\mid u\beta$. Furthermore, we have
$\sigma(S)=\frac{u\beta}{\alpha}R+h^{-1}\chi_{N}(\sigma)S+\beta(xR+yS)=\left(\frac{u\beta}{\alpha}+x\beta\right)R+(\chi_{N}(\sigma)+y\beta)S$
for some coefficients $x,y$.
@@
$\rho_{E_{2},N}(\sigma)=\begin{bmatrix}h+m\alpha&\frac{u\beta}{\alpha}+x\beta\\ n\alpha&h^{-1}\chi_{N}(\sigma)+y\beta\end{bmatrix}.$
Taking the determinant (viewed in $\mathbb{Z}/N\mathbb{Z}$), we obtain
$\chi_{N}(\sigma)=\chi_{N}(\sigma)+hy\beta+h^{-1}m\alpha\chi_{N}(\sigma)+my\alpha\beta-nu\beta.$
However, $u\beta$ is divisible by both $\alpha$ and $\beta$, so $u\beta=0$ as $\operatorname{lcm}(\alpha,\beta)=N$. Similarly, $my\alpha\beta=0$. Simplifying the equation with respect to these conclusions, we obtain
$hy\beta+h^{-1}m\alpha\chi_{N}(\sigma)=0.$
From here, we see that $\beta$ and $\alpha$ both divide $m\alpha$, so $m\alpha=0$, and $y\beta=0$ follows. In conclusion, we have
$\rho_{E_{2},N}(\sigma)=\begin{bmatrix}h&\frac{u\beta}{\alpha}+x\beta\\ n\alpha&h^{-1}\chi_{N}(\sigma)\end{bmatrix}.$
Finally, since $u\beta$ and $\alpha x\beta$ are both divisible by $N$, we see that $\frac{u\beta}{\alpha}$ and $x\beta$ are both divisible by $\frac{N}{\alpha}$, so the top right entry is always divisible by $\frac{N}{\alpha}$, which proves the lemma. $\Box$
In other words, if $E_{1}$ has an $N$-isogeny over $K$ for some $N$ and we apply some other cyclic isogeny $\phi:E_{1}\to E_{2}$, the $N$-isogeny turns into an $\alpha$-isogeny and an $\frac{N}{\alpha}$-isogeny on $E_{2}$, and those two are independent.
We also spell out the special case when $E_{1}$ has a point of order $N$ over $K$, not just an $N$-isogeny.
@@
Let $E_{1}/F$ be an elliptic curve with a point $P$ of order $N$ defined over $F$. Let $\phi:E_{1}\to E_{2}$ be a cyclic isogeny defined over $F$. Denote by $\alpha$ the order of $\phi(P)$ in $E_{2}$. Then the image of the mod $N$ Galois representation $\rho_{E_{1},N}(G_{F})$ is conjugate to a subgroup of the group of all matrices of the form
@@
The following result is well-known, but we reprove it for the sake of completion.
@@
[4, Proposition 3.3.] Let $E/\mathbb{Q}$ be a non-CM elliptic curve. Let $\ell$ be an odd prime. If an elliptic curve has no $\ell$-isogenies defined over $\mathbb{Q}$, then it also has no $\ell$-isogenies defined over $K$ for any odd-degree number field $K$, unless $\ell=7$ and $j(E)=2268945/128$, in which case there is a $7$-isogeny defined over the cubic field $\mathbb{Q}(\alpha)$, where $\alpha^{3}-5\alpha-5=0$.

### v2

@@ Abstract.
We determine all groups which occur as torsion subgroups of $\mathbb{Q}$-curves defined over number fields of degrees $3$, 55 and $7$. In particular, we prove that every torsion subgroup of a $\mathbb{Q}$-curve defined over a number field of degree $3,5$ or $7$ already occurs as a torsion subgroup of an elliptic curve with rational $j$-invariant. As the quadratic case has been solved by Le Fourn and Najman, and the case of extensions of prime degree greater than $7$ has been solved by Cremona and Najman, this paper completes the classification of torsion of $\mathbb{Q}$-curves over number fields of prime degree. To solve our problem, we study how images of Galois representations of elliptic curves change upon applying an isogeny. Since Cremona and Najman proved that $\mathbb{Q}$-curves over an odd degree number field $K$ are exactly those which are $K$-isogenous to an elliptic curve with rational $j$-invariant, we can rely on known results about images of Galois representations of rational elliptic curves.
As a corollary of our results, we establish that the torsion subgroup of an elliptic curve over a number field $K$ of prime degree which is isogenous to an elliptic curve with rational $j$-invariant is equal to the torsion subgroup of some elliptic curve defined over a degree $p$ number field with rational $j$-invariant.
@@ 1. Introduction
Let $E/K$ be an elliptic curve, where $K$ is a number field. By the Mordell–Weil theorem, the group $E(K)$ is finitely generated and hence of the form $T\times\mathbb{Z}^{r}$ for some nonnegative integer $r$, where $T$ is a finite Abelian group called the torsion subgroup of $E$.
A typical question one might ask is to determine all possible torsion subgroups of elliptic curves over a given number field. For elliptic curves over $\mathbb{Q}$, we have the following famous result by Mazur.
@@
$\mathbb{Z}/N_{1}\mathbb{Z}$ $\text{ with }N_{1}=1,\ldots,10,12,$
$\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}/2N_{2}\mathbb{Z}$ $\text{ with }N_{2}=1,\ldots,4.$
@@
In this article, we determine all possible torsion subgroups of $\mathbb{Q}$-curves over number fields of degree $3$, 55 and $7$. Let us first define $\mathbb{Q}$-curves. For a field $K$, we will denote its absolute Galois group $\operatorname{Gal}(\overline{K}/K)$ by $G_{K}$.
Definition 1.2.
We say an elliptic curve $E$ over a number field $K$ is a $\mathbb{Q}$-curve if it is $\overline{K}$-isogenous to all its $G_{\mathbb{Q}}$-conjugates.
For example, all elliptic curves isogenous to an elliptic curve with rational $j$-invariant are $\mathbb{Q}$-curves. Indeed, if $E/K$ is $\overline{K}$-isogenous to an elliptic curve $E^{\prime}$ with rational $j$-invariant, then there is a twist $E^{\prime\prime}$ of $E^{\prime}$ which is defined over $\mathbb{Q}$, and an isogeny $\varphi:E\to E^{\prime\prime}$. Then there is also an isogeny $\varphi^{\sigma}:E^{\sigma}\to(E^{\prime\prime})^{\sigma}$, but $(E^{\prime\prime})^{\sigma}=E^{\prime\prime}$, and the composition of $\varphi$ and the dual of $\varphi^{\sigma}$ is an isogeny from $E$ to $E^{\sigma}$.
Over number fields of odd degree, Cremona and Najman have proven that all non-CM $\mathbb{Q}$-curves arise in this way, with the isogeny even being defined over the base field of the $\mathbb{Q}$-curve.
@@
As any isogeny is a composition of a cyclic isogeny and the multiplication–by–$m$ map for some integer $m$, we may assume that the isogeny from the above proposition is cyclic.
@@
[15, Theorem 1.] Let $E/\mathbb{Q}$ be a rational elliptic curve, and let $K/\mathbb{Q}$ be a cubic extension. Then $E(K)_{tors}$ is isomorphic to one of the following $20$ groups:
$\mathbb{Z}/N_{1}\mathbb{Z}$ $\text{ with }N_{1}=1,\ldots,10,12,13,14,18,21,$
$\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}/2N_{2}\mathbb{Z}$ $\text{ with }N_{2}=1,2,3,4,7.$
@@
[7, Theorem 1] Let $E/\mathbb{Q}$ be a rational elliptic curve, and let $K/\mathbb{Q}$ be a quintic extension. Then $E(K)_{tors}$ is isomorphic to one of the following $17$ groups:
$\mathbb{Z}/N_{1}\mathbb{Z}$ $\text{ with }N_{1}=1,\ldots,10,11,12,25$
$\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}/2N_{2}\mathbb{Z}$ $\text{ with }N_{2}=1,2,3,4.$
@@
[6, follows from Proposition 7.7] Let $E/\mathbb{Q}$ be a rational elliptic curve, and let $K/\mathbb{Q}$ be a septic extension. Then $E(K)_{tors}$ is isomorphic to one of the $15$ groups from Theorem˜1.1.
@@
Gužvić proved in [9, Theorem 5.1.2] that elliptic curves with rational $j$-invariant over an odd prime degree number field have the same list of possible torsion subgroups as base changes of rational elliptic curves.
Over quadratic number fields, there is one exception. Namely, the group $\mathbb{Z}/13\mathbb{Z}$ appears among torsion subgroups of elliptic curves with rational $j$-invariant, and does not appear among torsion subgroups of rational elliptic curves. For example, consider the elliptic curve 100.1-e2 from the LMFDB.
@@
Let $K/\mathbb{Q}$ be a cubic extension and let $E/K$ be a $\mathbb{Q}$-curve. Then $E(K)_{tors}$ is isomorphic to one of the $20$ groups appearing in Theorem˜1.5.
@@
Let $K/\mathbb{Q}$ be a quintic extension and let $E/K$ be a $\mathbb{Q}$-curve. Then $E(K)_{tors}$ is isomorphic to one of the $17$ groups appearing in Theorem˜1.6.
@@
Let $K/\mathbb{Q}$ be a septic extension and let $E/K$ be a $\mathbb{Q}$-curve. Then $E(K)_{tors}$ is isomorphic to one of the $15$ groups appearing in Theorem˜1.7.
@@
$\mathbb{Z}/N\mathbb{Z},$ $\text{ for }N=13,14,15,16,18,$
$\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}/2N\mathbb{Z}$ $\text{ for }N=5,6,$
$\mathbb{Z}/3\mathbb{Z}\oplus\mathbb{Z}/3N\mathbb{Z}$ $\text{ for }N=1,2,$
@@
In [4, Theorem 1.3.], Cremona and Najman have shown that when $p>7$ is a prime and $K/\mathbb{Q}$ is an extension of degree $p$, every torsion subgroup $E(K)_{tors}$ of some $\mathbb{Q}$-curve $E/K$ is one of the subgroups already appearing in Theorem˜1.1.
@@
Under our definition, CM elliptic curves are also $\mathbb{Q}$-curves. In [3], Clark, Corn, Rice and Stankewicz classified torsion of CM elliptic curves over number fields of degree at most $13$. They proved that groups which occur as torsion subgroups of CM curves over cubic fields are those from Theorem˜1.1 along with $\mathbb{Z}/9\mathbb{Z}$ and $\mathbb{Z}/14\mathbb{Z}$. Over quintic fields, the only new group which appears is $\mathbb{Z}/11\mathbb{Z}$. Over septic fields, no new groups appear. In the rest of this article, we only focus on non-CM elliptic curves.
@@
The paper is structured as follows. In Section˜2, we list some general helpful results that will be used throughout the paper. In particular, we explicitly spell out the connection between the images of Galois representations of isogenous elliptic curves.
In Section˜3, we prove Theorem˜1.9, i.e. we solve the problem for cubic extensions. We do this case by case, eliminating each group which occurs for elliptic curves over cubic extensions and does not occur for rational elliptic curves over cubic extensions. Most of the techniques used in the cubic case are also used in the quintic and septic case, which are then discussed in Section˜4. In that case, we do not have a complete list of possible torsion subgroups, but we do have finitely many cases to check as there are finitely many degrees of cyclic isogenies of elliptic curves with rational $j$-invariant. Finally, in Section˜5, we conclude the proof of Theorem˜1.14.
Acknowledgements
The author thanks the anonymous referee for carefully reading the article, for numerous helpful comments, and for pointing out an error in a previous version, which has now been corrected.
@@ 2. General results
In this section, we recall some helpful results which will be used throughout the paper. Throughout the section, we will assume that all isogenies are cyclic.
@@
Let $E$ be an elliptic curve over a number field $K$. Let $E^{\prime}$ be a quadratic twist of $E$ by $d\in K^{\times}$. Let $\psi:G_{K}\to\{-1,1\}$ be the quadratic character defined by $\sigma(\sqrt{d})=\psi(\sqrt{d})\sqrt{d}$ for $\sigma\in G_{K}$. Denote by $\rho_{E,N}$ and $\rho_{E^{\prime},N}$ the mod $N$ Galois representations of $E$ and $E^{\prime}$ respectively. Then, for an appropriate choice of bases of $E[N]$ and $E^{\prime}[N]$, we have
@@
The proof can be found in [18, Lemma 5.1].
Since non-CM $\mathbb{Q}$-curves over odd degree number fields are isogenous to elliptic curves with rational $j$-invariants, it is very useful to describe what happens to the image of Galois representations after applying isogenies. We start with the following lemma which describes what happens with a point of prime order on the starting elliptic curve. We denote the cyclotomic character modulo $N$ by $\chi_{N}$.
@@
[2, Lemma 4.2.] Let $E_{1}/K$ be an elliptic curve with a point $P\in E_{1}(K)$ of prime order $p$ and let $\varphi:E_{1}\to E_{2}$ be an isogeny defined over $K$. Then either $E_{2}$ has a point of order $p$ defined over $K$ or for some basis of $E_{2}[p]$
$\rho_{E_{2},p}(\sigma)=\begin{bmatrix}\chi_{p}(\sigma)&y\\ 0&1\end{bmatrix},\ \text{for all }\sigma\in G_{K}.$
To avoid referring to bases, we may rephrase Lemma˜2.2 as follows.
@@
Let $E_{1}/F$ be an elliptic curve with a point $P\in E_{1}(K)$ of prime order $p$ and let $\varphi:E_{1}\to E_{2}$ be an isogeny defined over $K$. Then every $\alpha\in\rho_{E_{2},p}(G_{K})$ has eigenvalue $1$.
Combining Lemma˜2.1 and Lemma˜2.2, we obtain the following.
@@
Let $E$ be a $\mathbb{Q}$-curve over an odd degree number field $K$ with a point of order $p$ defined over $K$, where $p$ is a prime number. Let $\phi:E\to E^{\prime}$ be an isogeny, where $E^{\prime}$ is a curve with rational $j$-invariant, and let $E^{\prime\prime}$ be a quadratic twist of $E^{\prime}$ which is defined over $\mathbb{Q}$. Then $E^{\prime\prime}$ has a $p$-isogeny defined over $K$ and any element of $\rho_{E^{\prime\prime},p}(G_{K})$ has eigenvalue $1$ or 1-1.
The following proposition, which slightly generalizes [2, Proposition 4.4.], describes what happens in the composite case.
@@
Let $E_{1}/K$ be an elliptic curve with a point $P$ of order $N$ such that $\langle P\rangle$ is fixed by $G_{K}$. Denote by $H\leq(\mathbb{Z}/N\mathbb{Z})^{\times}$ the subgroup $\{k\in(\mathbb{Z}/N\mathbb{Z})^{\times}\mid\sigma(P)=kP\text{ for some }\sigma\in G_{K}\}$.
Let $\alpha$ be a divisor of $N$ and $\phi:E_{1}\to E_{2}$ the $\alpha$-isogeny defined over $K$ with kernel $\alpha\cdot P$. Then the image of the mod $N$ representation $\rho_{E_{2},N}(G_{K})$ is conjugate to a subgroup of the group of all matrices of the form
@@
Proof.   Let $\{P,Q\}$ be a basis for $E_{1}[N]$. We claim that $\phi(Q)$ has order $N$. Indeed, if $m\phi(Q)=0$ for $m<N$, then $mQ\in\ker\phi$. Since $\ker\phi$ is generated by a multiple of $P$, one has $mQ+dP=0$ for some integer $d$. However, $P$ and $Q$ are independent, so $m=d\alpha=0$.
Now let $P_{0}$ be a point on $E_{1}[N]$ such that $\frac{N}{\alpha}P_{0}=P$. Let $R=\phi(P_{0})$ and $S=\phi(Q)$. We claim that $\{R,S\}$ is a basis for $E_{2}[N]$. Both points are of order $N$, it remains to prove they are independent. If $aR+bS=0$, then $\phi(aP_{0}+bQ)=0$, so $aP_{0}+bQ=d\alpha P$ for some number $d$. It follows that $\frac{N}{\alpha}$ divides $a$ since $aP_{0}$ is of order $N$, and then we have a linear combination of $P$ and $Q$ being equal to 0. From here, we get $b=0$ and the independence is proven.
We now compute the action of $\sigma\in G_{K}$ on both elements of this basis. We know that
@@
Furthermore, $\sigma(S)=\sigma(\phi(Q))=\phi(\sigma(Q))$. If we set $\sigma(Q)=uP+vQ$, we get
$\sigma(S)=u\phi(P)+v\phi(Q)=\frac{N}{\alpha}uR+vS.$
Note that $v$ is the bottom right entry of $\rho_{E_{1},N}(\sigma)$, and $h$ is the top left entry of $\rho_{E_{1},N}(\sigma)$. Their product $hv$ is the determinant of $\rho_{E_{1},N}(\sigma)$ which is equal to the cyclotomic character $\chi_{N}$ by the Weil pairing. This does not depend on the elliptic curve.
@@
$\rho_{E_{2},N}(\sigma)=\begin{bmatrix}h+m\alpha&\frac{uN}{\alpha}\\ n\alpha&v\end{bmatrix}$.
Taking the determinant, we obtain
$\chi_{N}(\sigma)=(h+m\alpha)\cdot v.$
However, we know from before that $\chi_{N}(\sigma)=hv$. We get $m\alpha v=0$ and hence $m\alpha=0$. Thus, the top left entry of $\rho_{E_{2},N}(\sigma)$ is $h$. $\Box$
We also spell out the special case when $E_{1}$ has a point of order $N$ over $K$, not just an $N$-isogeny. This statement is in fact equivalent to [2, Proposition 4.4]
@@
Let $E_{1}/K$ be an elliptic curve with a point $P$ of order $N$ defined over $K$. Let $\alpha$ be a divisor of $N$ and $\phi:E_{1}\to E_{2}$ the $\alpha$-isogeny defined over $K$ with kernel $\alpha\cdot P$. Then the image of the mod $N$ representation $\rho_{E_{2},N}(G_{K})$ is conjugate to a subgroup of the group of all matrices of the form
@@
The following result is also well-known, but we reprove it for the sake of completion.
@@
[4, Proposition 3.3.] Let $\ell$ be an odd prime. If a non-CM elliptic curve has no $\ell$-isogenies defined over $\mathbb{Q}$, then it also has no $\ell$-isogenies defined over $K$ for any odd-degree number field $K$, unless $\ell=7$ and $j(E)=2268945/128$, in which case there is a $7$-isogeny defined over the cubic field $\mathbb{Q}(\alpha)$, where $\alpha^{3}-5\alpha-5=0$.
Throughout the article, we will often implicitly use the following basic fact.
Lemma 2.9.
Let $N$ be a positive integer and $E/K$ an elliptic curve. If $F/K$ is a finite extension of fields, then $\rho_{E,N}(G_{K}):\rho_{E,N}(G_{F})]$ divides $[F:K]$.
We remind the reader that possible torsion subgroups of CM elliptic curves are known by results from [3], so unless stated otherwise we will assume our elliptic curves are non-CM in what follows.
2.1. Isogeny graphs
By Proposition˜1.3, any non-CM $\mathbb{Q}$-curve over an odd degree number field $K$ is $K$-isogenous to an elliptic curve with rational $j$-invariant. In some of the proofs, we will use isogeny graphs of elliptic curves.
Let $K$ be a number field and $E/K$ a non-CM elliptic curve. Consider the graph whose vertices are all elliptic curves over $K$ which are $K$-isogenous to $E$, and for each prime number $p$, two vertices $E_{1}$ and $E_{2}$ are connected by an edge with label $p$ if and only if there is a $p$-isogeny from $E_{1}$ to $E_{2}$ which is defined over $K$.
This graph is called the $K$-isogeny graph of an elliptic curve $E/K$. We denote it by $\mathcal{G}(E/K)$. For each prime number $p$, the subgraph of $\mathcal{G}(E/K)$ induced by vertices which are $p^{k}$-isogenous to $E$ for some $k\geq 0$ is called the $p$-primary component $\mathcal{G}_{p}(E/K)$ of $\mathcal{G}(E/K)$. For non-CM elliptic curves, $\mathcal{G}_{p}(E/K)$ is a tree, for a cycle in the graph would imply the existence of an endomorphism which is not a multiplication-by-$m$ map.
It is well-known that for an odd prime $p$, any elliptic curve $E$ over a number field $K$ has either $0,1,2$ or $p+1$ isogenies of degree $p$ defined over $K$. For $p=2$, $E$ has either 0, 11 or $3$ isogenies of degree $2$. Furthermore, the following is true.
Lemma 2.10.
Let $p$ be a prime number. Let $E/K$ be an elliptic curve over a number field $K$ with $p+1$ isogenies of degree $p$ defined over $K$. Then $K$ contains the unique quadratic subfield of the cyclotomic field $\mathbb{Q}(\zeta_{p})$.
Proof.   If $E/K$ has all $p+1$ isogenies of degree $p$ defined over $K$, then $\rho_{E,p}(G_{K})$ consists only of scalar matrices, since the action of the Galois group fixes every subgroup of order $p$ of $E[p]\cong(\mathbb{Z}/p\mathbb{Z})^{2}$. However, the determinant of a scalar matrix is a square. From the Weil pairing, it follows that the unique quadratic subfield of $\mathbb{Q}(\zeta_{p})$ is contained inside $K$. $\Box$
Corollary 2.11.
If $p$ is an odd prime, $K/\mathbb{Q}$ is a number field of odd degree and $E/K$ is an elliptic curve, the $p$-primary component $\mathcal{G}_{p}(E/K)$ is a line graph.
Proof.   The graph $\mathcal{G}_{p}(E/K)$ is a tree and hence connected, and Lemma˜2.10 implies that any vertex has degree $0,1$ or $2$. The claim follows. $\Box$